Suppose that you are interested in determining whether the advice given by a physician during a routine physical examination is effective in encouraging patientsto stop smoking. In a study of current smokers, one group of patients was given a brief talk about the hazards of smoking and was encouraged to quit. A second group received no talk related to smoking. All patients were given a follow-up exam. In the sample of 150 patients who had received the talk, 25 reported that they had quit smoking. In the sample of 90 patients that hadn’t received a talk, 8 had ceased smoking.

a. Create a 2x2 table out of the data provided in the question.

b. Use Chi-square test to determine whether a difference exists in the proportion of smokers that had quit smoking between the two groups. Choose α = 0.05.
Answer #1

a)

below is contingency ta ble:

Smoking
quit not quit total
Talk received 25 125 150
not received 8 82 90
total 33 207 240

b)

null hypothesis: there is no  difference exists in the proportion of smokers that had quit smoking between the two groups

alternate hypothesis: there is a difference exists in the proportion of smokers that had quit smoking between the two groups

degree of freedom (row-1)*(column-1)=(2-1)*(2-1)=1

for 1 df and 0.05 level rejection region  \chi2 >3.841

applying chi square test:

Expected Ei=Σrow*Σcolumn/Σtotal quit not quit Total
received 20.625 129.375 150
not received 12.375 77.625 90
Total 33 207 240
chi square    χ2 =(Oi-Ei)2/Ei quit not quit Total
received 0.928 0.148 1.076
not received 1.547 0.247 1.793
Total 2.475 0.395 2.869

as test statsiic 2.869 is not in rejection region ; we can not rject null hypothesis

we do not have evidence at 0.05 level to conclude that there is a  difference exists in the proportion of smokers that had quit smoking between the two groups.